For 2% p.a. rate, the discount factor for perpetuity is ______, while the compound factor for perpetuity is ______.

  • A. 10; infinity
  • B. infinity; 10
  • C. 50; infinity
  • D. infinity; 50
Answer #1

Discount factor for Perpetuity = 1/(1+r) = 1/(1+0.02) = 50

Compound factor for perpetuity = (1+r)^n = (1+ 0.02)^infinite = infinity

Answer : C : 50, infinity (Thumbs up please)